A 15kg object falls a distance of 5.0 meters. How much work does gravity do on the object during this motion? Use the work-energy principle (from chapter 6- NO POTENTIAL ENERGY) to determine the speed of the object just as it reaches the ground. Assume there is no air resistance and the object starts from rest.
Answer #1

Work energy theoram

Mgh( work done by gravity) = 1/2*mv^2, mgh=15*9.81*5=735.75joule

V^2=2gh=

V=9.9m/s